OpenStudy (babyslapmafro):

When integrating rational functions by partial fractions, what do you do if the power of the numerator is the same as the denominator?

OpenStudy (babyslapmafro):

Do you solve normally, long-divide, or is there some other method?

OpenStudy (babyslapmafro):

\[\int\limits_{}^{}\frac{ }{ }\frac{ 3x^2-10 }{ x^2-4x+4 }dx\]

OpenStudy (anonymous):

i would factor the numerator and denominator first

OpenStudy (anonymous):

but yeah if the degrees are the same then long divide first

OpenStudy (babyslapmafro):

the numerator can't be factored, correct? \[\int\limits_{}^{}\frac{ 3x^2-10 }{ (x-2)(x-2) }dx\]

OpenStudy (babyslapmafro):

ok ill long-divide

OpenStudy (anonymous):

pauls online notes have a great tutorial on how to do partial fractions btw

OpenStudy (babyslapmafro):

can you long-divide with equal powers? this doesn't seem to work...

OpenStudy (babyslapmafro):

nvm i got it